# GPG: A Simple and Strong Reinforcement Learning Baseline for Model Reasoning

### v2

@@ Abstract
1 Introduction
Large Language Models (LLMs) have achieved substantial advancements, progressively narrowing the gap towards achieving Artificial General Intelligence (AGI) [38; 16; 2; 53; 57; 7]. Recently, LLMs exemplified by OpenAI o1 [38] and DeepSeek R1 [16], have adopted a strategy of generating intermediate reasoning steps before producing final answers. This approach has markedly improved their efficacy in domain-specific tasks [23; 13; 21; 26; 29; 22], such as mathematical reasoning. The remarkable success of this technology is mainly attributed to the Reinforcement Fine-Tuning (RFT) method [45; 46; 59; 28; 19]. Through the application of RFT, the models allocate additional time to “deliberate” prior to generating answers, thereby constructing intricate reasoning chains and subsequently enhancing overall model performance.
@@ 1 Introduction
  • We have analyzed and demonstrated the reward bias inherent in existing advantage functions and revealed the limitations of simplistic debiasing methods. Our exploration of the gradient estimate bias phenomenon has led us to propose a simple yet accurate gradient estimation (AGE) technique.
@@ 2.2 Group Policy Gradient
$F_{norm}$ is an optional normalization technique, which is commonly applied in conjunction with reward clipping to mitigate the impact of unexpected outlier values. One widely adopted practice is to employ standard variance normalization within a training batch [34; 45]. This approach helps stabilize the training process by reducing the variance of the reward signal, which is particularly important when dealing with environments where the magnitude of rewards can vary significantly, such as in different Atari games. By normalizing the reward signal, the model becomes less sensitive to extreme values, thereby improving the robustness and convergence of the training algorithm. However, in the reasoning tasks involving large models, the reward is typically well-defined and does not suffer from the same variance issues observed in other environments. As for the Math reasoning problem, it’s a common practice to award the right answer with 1.0 and the wrong answer with 0.0.
We utilize a basic Math Reasoning setting  [^1] https://github.com/huggingface/open-r1/blob/main/recipes/Qwen2.5-Math-7B/grpo/config_simple_rl.yaml of SimpleRL from open-r1 [11]. Specifically, We remove the format reward and only enable the accuracy reward for simplicity.
The critical component: $\hat{A}_{i,t}$, has been underexplored in prior reasoning research. This gap in the literature highlights the need for further investigation into the role and impact of $\hat{A}_{i,t}$ within reasoning tasks. There are two unresolved problems.
@@
The $\hat{A}_{i,t}$ should not introduce reward bias. Otherwise, bias deviates from the original problem formulation. GRPO [46] formulate it as $F_{norm}=\operatorname{std}\{R(o)\}$, which is essentially a function of $s_{t}$ in Equation 2 and explicitly introduces reward bias. Since we aim to solve the original problem, we don’t want to apply surrogate or bias. However, As shown in Table 1, if we remove this bias item, i.e. $F_{norm}=1$, it (43.9%) cannot clearly outperform GRPO (43.7%), which is opposite to the observation of a concurrent work Dr GRPO [31].
@@
It should be noted that the value $\alpha$ is not a constant and it varies across different sample batches. We also illustrate the $\alpha$ with different steps in Figure 2, which indicates the necessity of gradient correction. We As for multi-GPU training, to achieve more accurate gradient calculations, it is advisable to gather all non-zero gradient samples across all GPUs and compute the average gradient uniformly. This approach can be implemented through a custom gradient aggregation function, which leads to increased communication overhead. Instead, we derive another equivalent format, which doesn’t require extra cost and we provide the proof in section A. Therefore, given a batch sample, the objective can be written as
@@
In a scenario where we reject the $M$ examples and resample responses in a manner similar to the approach presented in a recent work [59] until $M$ equals 0, $\alpha$ is set to 1. However, this particular setting is not training efficient. The reason is that the training time is constrained by the worker that takes the longest to collect the desired examples. In contrast, our proposed method demonstrates superior efficiency. Moreover, it has the capability to automatically adjust the loss based on the performance on the sampled batch.
@@
Table 2: Comparison of reinforcement learning algorithms with various components.
RL algorithms vary significantly in their approaches to tackling variance and optimizing policies. Two key components in many RL algorithms are surrogate loss and policy constraints. Surrogate loss functions help stabilize training by providing a proxy objective that approximates the original goal but is easier to optimize. For example, in PPO [45], the surrogate loss is designed to prevent large updates to the policy, making training more stable by maintaining a balance between exploration and exploitation. (See Appendix C for detailed derivations.) However, surrogate loss functions can sometimes limit the flexibility of the policy updates, potentially leading to sub-optimal solutions. Policy constraints, such as KL divergence limits, ensure that the policy does not change too drastically between updates. These constraints are critical for maintaining stability in training, as they prevent erratic behavior and large deviations from known good policies. However, overly restrictive policy constraints can hinder the agent’s ability to explore new, potentially better strategies.
GRPO [46] introduces the use of reference models and policy constraints to stabilize training. The reference model serves as a foundation to provide additional guidance and regularization, ensuring more consistent and robust learning by minimizing variance in policy updates. Policy constraints further aid in preventing drastic policy changes, contributing to the stability of the learning process. Full derivations can be found in the Appendix C. Conversely, the introduction of GPG aims to offer greater flexibility by removing both the reference model and policy constraints. This approach fosters a more open exploration of the solution space, allowing the RL agent to develop novel strategies to achieve optimal results without the constraints imposed by prior models or divergence measures.
@@ 3.1 Experimental Setup
Dataset and Benchmarks In the unimodal scenario, we utilize datasets from multiple sources such as open-s1, open-rs [9], and MATH-lighteval [17] for training. Specifically, we train the DeepSeek-R1-Distill-Qwen-1.5B base model with the open-s1 dataset, resulting in the GPG-RS1 model. Similarly, training with the open-rs dataset produces the GPG-RS3 model. Furthermore, by employing the MATH-lighteval dataset on the Qwen2.5-Math-7B base model, we do ablations on this dataset. In order to make comparison with other recent work, we utilize the simplelr_qwen_level3to5 dateset [61]. These datasets encompass a wide range of problem types and difficulty levels. To assess the reasoning capabilities of the models, we employ five distinct mathematics-focused benchmark datasets: AIME24, MATH-500 [29; 17], AMC23, Minerva [27], and OlympiadBench [22].
@@
0:  $o~{}~{}[\text{shape: }(B,G,C,dim)]\leftarrow$ Model Output, $r\leftarrow$ Reward
1:  Calculate $\log\pi_{\theta}(o)[per\_token\_logps]$ based on $o$ and model $\pi_{\theta}$
2:  Calculate $\hat{A}[advantages]$ and $\alpha$ based on Eq.6 and Eq.7.
@@
GPG-7B 51.0 33.3 80.0 65.0 34.2 42.4
Table 4: The zero-shot pass@1 performance of the 7B models across five mathematical reasoning benchmarks. $\dagger$: reproduced results using the released code. $\ddagger$: results from [9].
@@ 3.2 Unimodal Task Evaluation
Mathematical Reasoning using 7B model. As illustrated in Table. 4, GPG-7B achieves an average score of 51.0% on the AIME24 dataset and 33.3% on MATH-500, outperforming other models such as Eurus-2-7B-PRIME [8], SimpleRL-Zero-7B [61] , and OpenReasoner-Zero-7B 8k [20]. Compared to the contemporary Oat-Zero-7B [31], GPG-7B consistently displays stronger performance across most benchmarks. Particularly in the Minerva, GPG-7B leads with a score of 34.2%, surpassing Oat-Zero-7B’s 29.0%.
@@ 4 Related Work
Large Model Reasoning. Recent advancements in both LLM and Multimodal Large Language Model (MLLM) have increasingly focused on enabling models to simulate human-like, stepwise reasoning processes. In the field of LLMs, researchers have pioneered methods such as Chain-of-Thought (CoT) prompting [38; 51; 24; 41; 60; 35; 58], Tree-of-Thought [56], Monte Carlo Tree Search [12; 54; 49], and the construction of complex SFT datasets [35], to enhance performance in reasoning tasks. Notably, approaches such as DeepSeek-R1 [16] have employed large-scale RL with format-specific and result-oriented reward functions, guiding LLMs toward self-emerging, human-like, complex CoT reasoning with significant performance improvements in challenging reasoning tasks. Meanwhile, MLLMs, which convert inputs from various modalities into a unified LLM vocabulary representation space for processing, have exhibited superior performance in vision understanding tasks [2; 53; 57; 7; 30; 5; 14; 37; 50]. Building on the advancements in LLM reasoning, there has been a collective effort within the research community to apply the DeepSeek-R1 methodology to MLLMs to enhance their visual reasoning capabilities, yielding remarkable progress [62; 32; 4; 64].
Reinforcement Learning. RL has driven significant progress in sequential decision-making, with policy gradient methods being fundamental to optimizing stochastic policies. The REINFORCE algorithm [52] established early principles for gradient-based policy updates in trajectory-driven tasks. However, its high variance has posed challenges for scalability. To address this, subsequent research has focused on stabilizing policy optimization processes. Trust Region Policy Optimization (TRPO) [43] introduced constrained updates via quadratic approximations to ensure monotonic improvement. This was further refined by PPO [45], which employed clipped objective functions to simplify the optimization process. Subsequent studies have sought to enhance the PPO alogrithm [6; 63] or elaborate on its implementation [10]. PPO has achieved widespread use in language model alignment and robotic control. However, the algorithm’s dependence on conservative policy updates or heuristic clipping thresholds can undermine its exploration potential in favor of stability, which poses a significant challenge in complex domains requiring dynamic strategy adaptation.
@@ Appendix B More Experiment Details
B.1 More Experiment Results
@@
B.2 Prompt and Reward Function
@@
B.3 More Experiment Settings
To evaluate the unimodal reasoning capabilities of our proposed method, we utilize two publicly available code repositories: Open-r1 [11] and Open-rs [9]. These repositories are selected due to their extensive coverage of various reasoning scenarios and their ability to present substantial challenges that effectively assess the reasoning capabilities of advanced models. The DeepSeek-R1-Distill-Qwen-1.5B model is trained for 100 and 50 global steps using the open-s1 and open-rs datasets, as reported in the repository [9], resulting in the GPG-RS1 and GPG-RS3 models, respectively. Subsequently, GPG-RS1 and GPG-RS3 are evaluated across five established benchmarks. Additionally, the Qwen2.5-Math-7B model in Table 1 is trained on 7,500 samples from the MATH-lighteval dataset, based on Open-r1 [11] code base.
@@ B.3 More Experiment Settings
Figure 4: Comparison of GPG(blue curves) and GRPO(gray curves) in terms of training loss, rewards and completion length. Experiments are based on DeepSeek-R1-Distill-Qwen-1.5B, same as Table. LABEL:tab:open-rs.

### v3

@@ 1 Introduction
Large Language Models (LLMs) have achieved substantial advancements, progressively narrowing the gap towards achieving Artificial General Intelligence (AGI) [38; 16; 2; 54; 58; 7]. Recently, LLMs exemplified by OpenAI o1 [38] and DeepSeek R1 [16], have adopted a strategy of generating intermediate reasoning steps before producing final answers. This approach has markedly improved their efficacy in domain-specific tasks [23; 13; 21; 26; 29; 22], such as mathematical reasoning. The remarkable success of this technology is mainly attributed to the Reinforcement Fine-Tuning (RFT) method [45; 46; 60; 28; 19]. Through the application of RFT, the models allocate additional time to “deliberate” prior to generating answers, thereby constructing intricate reasoning chains and subsequently enhancing overall model performance.
@@
  • We have analyzed and demonstrated the reward bias inherent in existing advantage functions and revealed the limitations of simplistic debiasing methods. Our exploration of the gradient estimate bias phenomenon has led us to propose a simple yet accurate gradient estimation (AGE) technique. To mitigate the potential issue of large variance in gradient estimation when the proportion of valid samples is excessively small, we introduce a simple thresholding mechanism to ensure a minimal partition of valid samples is maintained, followed by resampling.
@@ 2.2 Group Policy Gradient
$F_{norm}$ is an optional normalization technique, which is commonly applied in conjunction with reward clipping to mitigate the impact of unexpected outlier values. One widely adopted practice is to employ standard variance normalization within a training batch [34; 45]. This approach helps stabilize the training process by reducing the variance of the reward signal, which is particularly important when dealing with environments where the magnitude of rewards can vary significantly, such as in different Atari games. By normalizing the reward signal, the model becomes less sensitive to extreme values, thereby improving the robustness and convergence of the training algorithm. However, in the reasoning tasks involving large models, the reward is typically well-defined and does not suffer from the same variance issues observed in other environments. As for the Math reasoning problem, it is a common practice to award the right answer with 1.0 and the wrong answer with 0.0.
We utilize a basic Math Reasoning setting  [^1] https://github.com/huggingface/open-r1/blob/main/recipes/Qwen2.5-Math-7B/grpo/config_simple_rl.yaml of SimpleRL from open-r1 [11], using only the MATH-lighteval dataset to facilitate rapid experimental validation. Specifically, we remove the format reward and only enable the accuracy reward for simplicity.
@@
The critical component: $\hat{A}_{i,t}$, has been underexplored in prior research in reasoning. This gap in the literature highlights the need for further investigation of the role and impact of $\hat{A}_{i,t}$ within reasoning tasks. There are two unresolved problems.
The $\hat{A}_{i,t}$ should not introduce reward bias. Otherwise, bias deviates from the original problem formulation. GRPO [46] formulates it as $F_{norm}=\operatorname{std}\{R(o)\}$, which is essentially a function of $s_{t}$ in Equation 2 and explicitly introduces the reward bias. Since we aim to solve the original problem, we don’t want to apply surrogate or bias. However, As shown in Table 1, if we remove this bias item, i.e. $F_{norm}=1$, it (43.9%) cannot clearly outperform GRPO (43.7%), which is opposite to the observation of a concurrent work Dr. GRPO [31].
@@
It should be noted that the value $\alpha$ is not a constant and it varies across different sample batches. We also illustrate $\alpha$ with different steps in Figure 2, which indicates the necessity of gradient correction. We As for multi-GPU training, to achieve more accurate gradient calculations, it is advisable to gather all non-zero gradient samples across all GPUs and compute the average gradient uniformly. This approach can be implemented through a custom gradient aggregation function, which leads to increased communication overhead. Instead, we derive another equivalent format, which doesn’t require extra cost and we provide the proof in section A. Therefore, given a batch sample, the objective can be written as
@@
In a scenario where we reject the $M$ examples and resample responses in a manner similar to the approach presented in a recent work [60] until $M$ equals 0, $\alpha$ is set to 1. However, this particular setting is not training efficient. The reason is that the training time is constrained by the worker that takes the longest to collect the desired examples. In contrast, our proposed method demonstrates superior efficiency. Moreover, it has the capability to automatically adjust the loss based on the performance of the sample batch.
@@
Table 2: Comparison of reinforcement learning algorithms (in reasoning) with various components.
Thresholding minimal partition of valid samples and resampling to reduce variance. While our approach provides an unbiased estimation of the gradient, it may encounter issues with high variance when the proportion of valid samples is excessively low. To mitigate this, we introduce a threshold $\beta_{th}=\frac{1}{\alpha_{th}}$ for the proportion of valid samples. When this proportion falls below the given value, we accumulate the valid samples into the resampled subsequent batch until the proportion exceeds the threshold. This strategy effectively reduces the variance of the gradient estimation, thereby enhancing the stability and convergence rate of the model training process. It is important to highlight that this strategy can further enhance the performance, as demonstrated in Table 1.
RL algorithms vary significantly in their approaches to tackling variance and optimizing policies. Two key components in many RL algorithms are surrogate loss and policy constraints. We summarize the main comparisons among various frameworks in Table 2. Our method stands out by preserving the simplest form, which not only ensures ease of implementation but also maintains high efficiency and effectiveness.
@@ 3.1 Experimental Setup
Dataset and Benchmarks In the unimodal scenario, we utilize datasets from multiple sources such as open-s1, open-rs [9], and MATH-lighteval [17] for training. Specifically, we train the DeepSeek-R1-Distill-Qwen-1.5B base model with the open-s1 dataset, resulting in the GPG-RS1 model. Similarly, training with the open-rs dataset produces the GPG-RS3 model. Furthermore, by employing the MATH-lighteval dataset on the Qwen2.5-Math-7B base model, we do ablations on this dataset. To compare the overall performance on the 7B model, we utilize the dataset from [60] and the detailed setting is shown in section B.1.
These datasets encompass a wide range of problem types and difficulty levels. To assess the reasoning capabilities of the models, we employ five distinct mathematics-focused benchmark datasets: AIME24, MATH-500 [29; 17], AMC23, Minerva [27], and OlympiadBench [22].
@@
0:  $o~{}~{}[\text{shape: }(B,G,C,dim)]\leftarrow$ Model Output, $r\leftarrow$ Reward, $\beta_{th}$
1:  Collecting samples and calculate $\hat{A}[advantages]$ and $\alpha$ based on Eq.6 and Eq.7 until $\alpha<\frac{1}{\beta_{th}}$
2:  Calculate $\log\pi_{\theta}(o)[per\_token\_logps]$ based on $o$ and model $\pi_{\theta}$
@@
GPG-7B 57.7 36.7 84.6 82.5 39.0 45.8
Table 4: The zero-shot pass@1 performance of the 7B models across five mathematical reasoning benchmarks. $\dagger$: reproduced results using the released code. $\ddagger$: results from [9], : results from [31].
@@ 3.2 Unimodal Task Evaluation
Mathematical Reasoning using 7B model. As illustrated in Table. 4, GPG-7B achieves an average score of 57.7% and outperforms other baselines with clear margins.
@@ 4 Related Work
Large Model Reasoning. Recent advancements in both LLM and Multimodal Large Language Model (MLLM) have increasingly focused on enabling models to simulate human-like, stepwise reasoning processes. In the field of LLMs, researchers have pioneered methods such as Chain-of-Thought (CoT) prompting [38; 52; 24; 41; 61; 35; 59], Tree-of-Thought [57], Monte Carlo Tree Search [12; 55; 50], and the construction of complex SFT datasets [35], to enhance performance in reasoning tasks. Notably, approaches such as DeepSeek-R1 [16] have employed large-scale RL with format-specific and result-oriented reward functions, guiding LLMs toward self-emerging, human-like, complex CoT reasoning with significant performance improvements in challenging reasoning tasks. Meanwhile, MLLMs, which convert inputs from various modalities into a unified LLM vocabulary representation space for processing, have exhibited superior performance in vision understanding tasks [2; 54; 58; 7; 30; 5; 14; 37; 51]. Building on the advancements in LLM reasoning, there has been a collective effort within the research community to apply the DeepSeek-R1 methodology to MLLMs to enhance their visual reasoning capabilities, yielding remarkable progress [63; 32; 4; 65].
Reinforcement Learning. RL has driven significant progress in sequential decision-making, with policy gradient methods being fundamental to optimizing stochastic policies. The REINFORCE algorithm [53] established early principles for gradient-based policy updates in trajectory-driven tasks. However, its high variance has posed challenges for scalability. To address this, subsequent research has focused on stabilizing policy optimization processes. Trust Region Policy Optimization (TRPO) [43] introduced constrained updates via quadratic approximations to ensure monotonic improvement. This was further refined by PPO [45], which employed clipped objective functions to simplify the optimization process. Subsequent studies have sought to enhance the PPO alogrithm [6; 64] or elaborate on its implementation [10]. PPO has achieved widespread use in language model alignment and robotic control. However, the algorithm’s dependence on conservative policy updates or heuristic clipping thresholds can undermine its exploration potential in favor of stability, which poses a significant challenge in complex domains requiring dynamic strategy adaptation.
@@ Appendix B More Experiment Details
B.1 Experiment Settings
Training setting on 7B based on dataset from [60]. We employ the VERL framework [47] with a global batch size of 144 prompts. For each prompt, we generate 8 responses and use only accuracy-based rewards. Our implementation strictly follows Algorithm 1. We optimize the network using the AdamW optimizer with a constant learning rate of 1$\times$ 10-6 and a weight decay of 0.1. The threshold value $\beta_{th}$ is set to 0.6. We trained the model for 1100 steps utilizing 48 NPUs sourced from China.
B.2 More Ablation Experiment Results
$\beta_{th}$ Average AIME24 MATH-500 AMC23 Minerva OlympiadBench
0.6 48.3 30.0 76.2 62.5 34.2 39.0
0.8 48.6 33.3 73.6 67.5 29.4 39.3
Table 10: Ablation on different $\beta_{th}$ using Qwen2.5 Math 7B.
@@ B.2 More Ablation Experiment Results
B.3 Prompt and Reward Function
@@
B.4 More Experiment Settings
To evaluate the unimodal reasoning capabilities of our proposed method, we utilize two publicly available code repositories: Open-r1 [11] and Open-rs [9]. These repositories are selected due to their extensive coverage of various reasoning scenarios and their ability to present substantial challenges that effectively assess the reasoning capabilities of advanced models. The DeepSeek-R1-Distill-Qwen-1.5B model is trained for 100 and 50 global steps using the open-s1 and open-rs datasets, as reported in the repository [9], resulting in the GPG-RS1 and GPG-RS3 models, respectively.
@@ B.4 More Experiment Settings
Figure 4: Comparison of GPG(blue curves) and GRPO(gray curves) in terms of training loss, rewards and completion length. Experiments are based on DeepSeek-R1-Distill-Qwen-1.5B, same as Table. 3.